Under a pressure of 3000N/m2, a gas has a volume of 250cm3. What will be its volume if the pressure is changed to 100mmHg at the same temperature?
[ 760mmHg = 101325N/m2]
1
Expert's answer
2020-04-27T01:46:29-0400

This task is based on Boyle's Law: "PV=const"

So "P_1V_1=P_2V_2" , P1 and P2 - initial and final pressure of a gas respectively, V1 and V2 - initial and final volume of a gas respectively.

Express V2: "P_1V_1\/P_2=V_2"

Pressure must be converted to the same units, use proportion:

760mmHg=101325N/m2

100mmHg=x (P2)

"x=100mmHg*101325N\/m^2\/760mmHg=13332N\/m^2"

V2=3000N/m2*250cm3/13332N/m2=56,3cm3

Answer: 56,3cm3.